OpenStudy (anonymous):

can any body check me work on chemistry I really appreciated.thank you I have the Majority of it .

OpenStudy (aaronq):

are you going to post it?

OpenStudy (anonymous):

yes

OpenStudy (aaronq):

can you post it as a pdf, I'm not a fan of word documents.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I don't know how to do it

OpenStudy (aaronq):

open it i word and click "save as", select the pdf option.

OpenStudy (anonymous):

OK I figure it out here it is.

OpenStudy (anonymous):

if you can not help me on this can you find a way i am not good in chemistry. thank you a I really appreciated.

OpenStudy (aaronq):

so you need help with: "5. Determine the percent yield of NaCl in your experiment, showing all work neatly in the space below. 6. If you had not heated the product long enough to remove all of the water, explain in detail how that would have specifically affected your calculated actual yield and percent yield." ?

OpenStudy (anonymous):

basically yes, I also want to know if number 4 is correct i feel that i am missing something.

OpenStudy (aaronq):

so for 5. you need to use this equation: \(percent~yield=\dfrac{experimental~yield}{theoretical~yield}*100\%\) But you need the results from q4 and q3. Question 4. The experimental yield is given in the lab: salt=(salt and dish) - (empty dish)=31.52 (g)-24.35 (g)=7.17 g Question 3. The theoretical you have to find using simple stoichiometry: \(NaHCO_3 + HCl → CO_2 + H_2O + NaCl\) start by converting the mass of NaHCO3 to moles. Use: \(moles=\dfrac{mass}{molar~mass}\) after you need to set up a ratio to find the moles of NaCl produced: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) From here you can isolate what you need. For example: if you have 2 moles of B, how many moles of C can you produce? solve algebraically: \(\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_C=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\) -------------------------------------------------------- To convert moles of NaCl to mass, use: \(moles=\dfrac{mass}{molar~mass}\) again. Plug back in the first formula.

OpenStudy (aaronq):

Q6. your experimental (actual) result would've appeared to be larger (i.e. more mass produced) because the water would add weight. This would make the percent yield higher (refer to formula from question 5), and the reaction would appear to be more efficient than it actually was.

OpenStudy (anonymous):

thank you for your help.

OpenStudy (aaronq):

no problem